OpenStudy (anonymous):

Give an example of a relation that is NOT a function and explain why it is not a function

OpenStudy (anonymous):

Anything that returns more than one value for a single input is not a function.

OpenStudy (anonymous):

One good example of this are circles.

OpenStudy (anonymous):

x^2+y^2=r^2 y= +- sqrt(r^2-x^2)